OpenStudy (anonymous):

A van traveling down a slope with a uniform acceleration of 2.15 meters/second2 attains a speed of 20.00 meters/second after 7.00 seconds. What is the initial velocity of the van? 0.00 meters/second 3.65 meters/second 4.95 meters/second 35.05 meters/second 120.75 meters/second

OpenStudy (shamim):

v=u+at

OpenStudy (shamim):

here a=2.15m/s^2 v=20m/s t=7s

OpenStudy (shamim):

just put these data into my equation

OpenStudy (shamim):

nd feel free to ask any question

OpenStudy (anonymous):

So it's c

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

Hello can you please help me with the answer

OpenStudy (shamim):

ok

OpenStudy (shamim):

ya u r correct

OpenStudy (shamim):

@Mahnoor